A gasoline pump delivers 4 2/5 gallons of gas per minute. How many minutes will it take to fill a gas tank that holds 16 1/2 gallons?
Answer by ptaylor(2198)

OK
Let x=number of minutes it takes to fill the tank
Then our equation to solve is:
(4 2/5 gal/min)*(x min)=16 1/2 gal
4 2/5=22/5=44/10
and
16 1/2=33/2=165/10
sooooo
(44/10)*x=165/10 multiply each side by 10
44x=165
x=3 3/4 min----------- time required to fill the tank
CK
(4 2/5)*(3 3/4)=
(22/5)*(15/4)=
(22)*(3/4)=
(11)*(3/2)=
33/2=16 1/2
16 1/2= 16 1/2
Hope this helps---ptaylor